The population N(t) of a species of micro-organism in a laboratory setting at any time t is established to vary under the influence of a certain chemical at a rate given by dN/dt=t-2et show that N(t)=t-2et+c.hence if N(0)=400 determine the population when t=5


1
Expert's answer
2021-07-08T09:20:53-0400

Given -


"\dfrac{dN}{dt}=" "t-2e^{t}"



Then , "dN=(t-2e^{t})dt"


integrate both sides we get ,



"\int dN=\int(t-2e^{t})dt"



"N(t)=\dfrac{t^{2}}{2}-2e^{t}+c"



given -


"N(0)=400"



"=400=\dfrac{0^{2}}{2}-2e^{0}+c" "\implies" "c=402"


"N(t)=" "\dfrac{t^{2}}{2}-2e^{t}+402"


now putting t=5 we get ,


"N(5)=""\dfrac{5^{2}}{2}-2e^{5}+402=118"